# LiteGS: A High-performance Framework to Train 3DGS in Subminutes via System and Algorithm Codesign

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as promising alternative in 3D representation. However, it still suffers from high training cost. This paper introduces LiteGS, a high performance framework that systematically optimizes the 3DGS training pipeline from multiple aspects. At the low-level computation layer, we design a “warp-based raster” associated with two hardware-aware optimizations to significantly reduce gradient reduction overhead. At the mid-level data management layer, we introduce dynamic spatial sorting based on Morton coding to enable a performant “Cluster-Cull-Compact” pipeline and improve data locality, therefore reducing cache misses. At the top-level algorithm layer, we establish a new robust densification criterion based on the variance of the opacity gradient, paired with a more stable opacity control mechanism, to achieve more precise parameter growth. Experimental results demonstrate that LiteGS accelerates the original 3DGS training by up to 13.4x with comparable or superior quality and surpasses the current SOTA in lightweight models by up to 1.4x speedup. For high-quality reconstruction tasks, LiteGS sets a new accuracy record and decreases the training time by an order of magnitude.
(b) PSNR versus training time. LiteGS converges to high-quality results over 10× faster than prior work, demonstrating strong advantages in both compactness and training efficiency.
Figure 1: Reconstruction Quality under Varying Model Budgets on the MipNeRF360 Garden Scene.
Introduction
As an emerging technique for scene representation and novel view synthesis, 3DGS has achieved remarkable breakthroughs in computer graphics and computer vision due to its high rendering quality and rapid rendering speed (Kerbl et al. 2023). This technique represents 3D scenes using millions of anisotropic 3D Gaussian primitives to obtain photorealistic rendering results, demonstrating immense potential in fields such as autonomous driving, virtual reality, and digital twins (Lin et al. 2024; Liu et al. 2024; Tu et al. 2025; Kung et al. 2025). Although 3DGS rendering is very fast, the training process could take tens of minutes or even hours, which limits its application. Existing efforts have attempted to reduce the number of Gausians or simplify certain computation to accelerate 3DGS training, but are still unable to mitigate the main performance bottlenecks (Mallick and Banerjee 2024; Durvasula and Chai 2024; Liao et al. 2025).
@@ Introduction
The system and algorithm co-optimization allows LiteGS to obtain significant superiority in both training efficiency and reconstruction quality, setting a new performance benchmark for the field. LiteGS can achieve the same quality as 3DGS-MCMC (quality SOTA) (Kheradmand et al. 2024) using less than half the number of parameters and up to 10.78x speedup. When the number of parameters increases to the same level, LiteGS surpasses SOTA solutions by 0.2-0.4 dB in PSNR, while shortening the training time to 6-8 minutes. For lightweight models, LiteGS requires only about 10% of the training time and 20% of the parameters to achieve the same quality as original 3DGS (Kerbl et al. 2023). Compared to other cutting-edge acceleration works in lightweight models like Mini-splatting v2 (Fang and Wang 2024b), LiteGS also demonstrates a speedup up to 1.4x. Comprehensive ablation studies further quantify and confirm the effectiveness and necessity of each designed module in our framework.
In summary, the main contributions of this paper are as follows:
@@
  • We identify the major bottleneck in the existing 3DGS training pipeline, and propose an array of techniques, e.g., warp-based raterization, Cluter-Cull-Compact, and novel densification strategies, to eliminate these bottlenecks.
@@
  • We propose a high-performance framework, LiteGS, to systematically accelerate 3DGS training through system and algorithm codesign. We have open-sourced the framework to the community [^1] https://github.com/MooreThreads/LiteGS.
@@
  • We perform extensive experiments and ablation studies to evaluate the effectiveness of LiteGS. Compared to the original 3DGS framework, it reduces the training time by up to 13.4x while still retaining comparable quality. In high-quality setting, it accelerates the SOTA acceleration solution by up to 10.7x with even slightly better quality.
Relate Work
Acceleration
@@
Although 3DGS is already faster to optimize than classic NeRF, the training still takes tens of minutes or even hours. A key focus in recent 3DGS research has been on removing redundant computations to speed up training. Several works reduce the per-pixel Gaussian overhead by culling or skipping negligible contributions (Liu et al. 2025; Hanson et al. 2025a; Feng et al. 2025). Hou et al. (2025) approximate the alpha blending with a commutative weighted sum, removing the need for per-frame sorting and modestly improving performance on mobile GPUs. Another line of work is pruning redundant primitives. LightGaussian (Fan et al. 2024) proposes a heuristic pruning pipeline. PUP 3D-GS (Hanson et al. 2025b) uses a Hessian-based sensitivity metric to remove  90% of Gaussians from a pretrained model while preserving detail. Other works try to improve the GPU utilizations of 3DGS training. To reduce atomic add operations in the gradient accumulation step, DISTWAR introduced an optimized reduction primitive that leverages warp-level parallelism(Durvasula and Chai 2024). To remove atomic conflicts on the same tile, Taming 3DGS proposed a ’per-splat parallelization’ approach for backpropagation (Mallick and Banerjee 2024). TC-GS uses Tensor Core units for splatting operations, formulating alpha-blending calculations as matrix multiplies so that otherwise underutilized hardware can accelerate them (Liao et al. 2025).
Densification Strategies
@@
A key to 3DGS’s success is its ability to refine the point cloud during training via Adaptive Density Control (ADC) - i.e. adding Gaussians in undersampled regions and removing those with negligible opacity. However, the original ADC heuristic has limitations that often produce an explosion of Gaussians with diminishing importance. Recent works have therefore revised the densification logic to be more principled and efficient. Bulò, Porzi, and Kontschieder (2024) propose an error-driven densification scheme: instead of using just gradient magnitude, they compute an auxiliary per-pixel error map and spawn new Gaussians where the reconstruction error remains high. Deng et al. (2024) tackle the root cause of Gaussian overshoot by replacing the clone-and-split step with a single long-axis split that avoids overlapping the child Gaussians. Their efficient density control (EDC) method achieves the same or better fidelity using only  1/3 of the Gaussians, by preventing the creation of redundant, overlapping splats. Other strategies guide densification using additional cues. Kim, Lee, and Lee (2024) incorporate color residuals (via spherical harmonics gradients) into the densification criterion, which helps target new Gaussians to areas of color discrepancy rather than just geometric edges. Taming 3DGS (Mallick et al. 2024) uses a score-based selection that adds only the Gaussians which maximally improve the scene’s SSIM/PSNR, and it grows the model to a predetermined budget (rather than indefinitely). By steering new Gaussians to where they raise fidelity the most and stopping when the budget is reached, this method obtains high reconstruction quality with 4-5× fewer Gaussians than the naive ADC (Mallick and Banerjee 2024).
Methodology
@@
where $s(i)$ denotes the index mapping function for depth-ordered primitives in camera space, corresponding to the $i$-th sorted primitive in the Gaussian scene representation $\mathbb{G}$, $T$ is the transmittance $T_{s(i),\pi_{j}}(x,y)=T_{s(i-1),\pi_{j}}(x,y)\times(1-\alpha_{s(i),\pi_{j}}(x,y))$, $N^{\prime}$ represents the number of primitives visible within the view frustum, $\alpha$ is the opacity term $\alpha_{i,\pi_{j}}(x,y)=o_{i}\times G(x-x_{i,\pi_{j}}^{screen},y-y_{i,\pi_{j}}^{screen},\Sigma_{i,\pi_{j}}^{screen})$ and $G$ is defined as,
$G(\Delta x,\Delta y,\Sigma)=e^{-0.5\times(\Delta x,\Delta y)\Sigma(\Delta x,\Delta y)^{T}}$ (2)
We use $P_{j}$ to represent the $j$-th rendering tile, and the gradient of $c^{r}_{s(i)}$ is computed by Eq. 3.
$\frac{\partial Loss_{\pi}}{\partial c^{r}_{s(i)}}=\sum_{j=0}^{M}\sum_{(x,y)\in P_{j}}\frac{\partial Loss_{\pi}}{\partial I_{r}(x,y)}\times T_{s(i),\pi}(x,y)\times\alpha_{s(i),\pi}(x,y)$ (3)
Warp-based Raster
@@
To remarkably eliminate the overhead, we propose a novel warp-based rasterization that only requires a single warp reduction. The key idea is to assign only one warp to each render tile, and let each thread within the warp process multiple pixels through iterative loops. This design first performs an intra-thread reduction operation to accumulate the gradients generated by each pixel within a thread, and then applies a single warp-level reduction to aggregate the gradients for the entire tile. By shrinking the number of expensive warp reductions required per tile to just one, we substantially improve the performance of gradient reduction.
However, the design does not come for free. It possibly leads to sharply increased register pressure because the approach tends to assign more pixels to a signle thread, which inevitably needs to allocate more temporary variables (such as color and opacity during alpha blending) in the local register. Excessive register occupation causes register spilling, which in turn leads to notable GPU under-utilization. We propose two techniques to ease the register pressure.
Scanline Algorithm
The first technique, inspired by the scanline algorithm(Foley et al. 1996; Shirley, Marschner et al. 2015) used in software rasterization, reforms the computation of the 2D Gaussian function to maximize data reuse hence diminish register allocation in each thread. This idea lies in the observation that there is a considerable amount of redundant computations during processing multiple pixels for the same Gaussian primitive.
We assume $\Sigma^{\text{screen}}=\begin{pmatrix}a&b\\ b&c\end{pmatrix}$. Along any scanline (e.g., the y-axis), the exponent of the Gaussian function (Eq. 2) can be decomposed into a quadratic polynomial of the pixel offset $j$: $\text{Basic}+\text{Linear}\cdot j+\text{Quad}\cdot j^{2}$. As shown in Algorithm 1, the coefficients-Basic, Linear, and Quad-are derived from the covariance matrix and the initial pixel’s offset. This structure allows for significant optimization: the coefficients are computed only once per scanline. The values for all subsequent pixels are then found efficiently through a few incremental multiply-add operations.
Algorithm 1 Scanline algorithm
[…]
Parameter: $length$
Output: $out\_color\in\mathbb{R}^{length\times 3}$
[…]
2:for $i\leftarrow 0$ to $N$ do
3:  $\Delta x=x_{i}^{screen}-x$
4:  $\Delta y=y_{i}^{screen}-y$
5:  $basic=-0.5\times(\Delta x,\Delta y)\times\Sigma_{i}^{screen}\times(\Delta x,\Delta y)^{T}$
6:  $linear=b\Delta x+c\Delta y$
7:  $quad=-0.5d$
8:  for $j\leftarrow 0$ to $length$ do
9:   $g=exp(basic+linear*j+quad*j^{2})$
10:   $alpha=g*o^{i}$
11:   $out\_color[j]+=T[j]*alpha*c_{i}$
12:   $T[j]*=(1-alpha)$
13:  end for
14:end for
[…]
Strategic Mixed-Precision Computing
@@
These synergistic optimizations help us effectively tune the best performance given the hardware resource constraints. For example, based on the optimized per-pixel register usage, we discovered that each thread processes 4 pixels is the optimal configuration. Using the “one warp per tile” principle, we identified that tile size with 4×32=1284$\times$ 32=1284 × 32 = 128 pixels yields the best performance. This size fits the major NVIDIA graphic GPUs since they have the same number of registers.
Cluster-Cull-Compact
@@
To address these challenges, we adapt the established “Cluster-Cull-Compact” paradigm from the field of real-time rendering (Laine, Karras, and Aila 2011; Hapala, Bittner, and Havran 2011), tailoring it for the dynamic training process of 3DGS. This pipeline consists of three stages: 1) grouping primitives into spatially contiguous clusters; 2) performing culling at the cluster granularity; and 3) compacting the data of visible clusters into a continuous memory block.
However, traditional clustering methods like Bounding Volume Hierarchies (BVH) or graph partitioning(Karis 2021) have high computational complexity for a training loop. The cornerstone of our approach is the utilization of Morton coding (Z-order curve) to achieve high-performance locality-aware sorting(Morton 1966). By mapping 3D coordinates to a 1D sequence that preserves locality, Morton coding enables us to efficiently enforce spatial coherence at a minimal cost and facilitate lightweight clustering operations.
Morton Codes
Our pipeline begins by normalizing the Gaussian coordinates $p_{i}=(x_{i},y_{i},z_{i})$ into the unit cube $[0,1]^{3}$. To generate a 64-bit Morton code, the normalized coordinates are scaled by $2^{22}-1$ and converted to integers. The 64-bit Morton code is then generated by interleaving the bits of the integer coordinates: $\text{Morton}(x,y,z)=\bigoplus_{k=0}^{20}\left(x_{k}\ll(2k)\,|\,y_{k}\ll(2k+1)\,|\,z_{k}\ll(2k+2)\right)$, where $\oplus$ denotes bit concatenation, and $x_{k}$ represents the $k$-th bit of the quantized coordinates. Fig. 2 illustrates how Morton codes map 3D points to a 1D sequence.

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as promising alternative in 3D representation. However, it still suffers from high training cost. This paper introduces LiteGS, a high performance framework that systematically optimizes the 3DGS training pipeline from multiple aspects. At the low-level computation layer, we design a “warp-based raster” associated with two hardware-aware optimizations to significantly reduce gradient reduction overhead. At the mid-level data management layer, we introduce dynamic spatial sorting based on Morton coding to enable a performant “Cluster-Cull-Compact” pipeline and improve data locality, therefore reducing cache misses. At the top-level algorithm layer, we establish a new robust densification criterion based on the variance of the opacity gradient, paired with a more stable opacity control mechanism, to achieve more precise parameter growth. Experimental results demonstrate that LiteGS accelerates the original 3DGS training by up to 13.413.4x with comparable or superior quality and surpasses the current SOTA in lightweight models by up to 1.51.5x speedup. For high-quality reconstruction tasks, LiteGS sets a new accuracy record and decreases the training time by an order of magnitude.
1 Introduction
As an emerging technique for scene representation and novel view synthesis, 3DGS has achieved remarkable breakthroughs in computer graphics and computer vision due to its high rendering quality and rapid rendering speed (Kerbl et al., 2023). This technique represents 3D scenes using millions of anisotropic 3D Gaussian primitives to obtain photorealistic rendering results, demonstrating immense potential in fields such as autonomous driving, virtual reality, and digital twins (Lin et al., 2024; Liu et al., 2024; Tu et al., 2025; Kung et al., 2025). Although 3DGS rendering is very fast, the training process could take tens of minutes or even hours, which limits its application. Existing efforts have attempted to reduce the number of Gausians or simplify certain computation to accelerate 3DGS training, but are still unable to mitigate the main performance bottlenecks (Mallick et al., 2024; Durvasula & Chai, 2024; Liao et al., 2025).
@@ 1 Introduction
The system and algorithm co-optimization allows LiteGS to obtain significant superiority in both training efficiency and reconstruction quality, setting a new performance benchmark for the field. LiteGS can achieve the same quality as 3DGS-MCMC (quality SOTA) (Kheradmand et al., 2024) with up to 10.810.8x speedup using less than half the number of parameters. However, using the same amount of parameters, LiteGS surpasses the quality SOTA solutions by 0.20.2-$0.4$ dB in PSNR and shortens the training time by 3.83.8x to 77x. For lightweight models, LiteGS requires only about $10\%$ training time and $20\%$ parameters to produce the same quality as the original 3DGS (Kerbl et al., 2023). Compared to other cutting-edge acceleration works in lightweight models such as Mini-splatting v2 (Fang & Wang, 2024b), LiteGS also reduces the training cost by up to 1.51.5x. Comprehensive ablation studies further quantify and confirm the effectiveness and necessity of each technique proposed in our framework.
In summary, the main contributions of this paper are as follows.
@@
  • We identify the major bottlenecks in the existing 3DGS training pipeline, and propose an array of techniques, e.g., warp-based rasterization, Cluster-Cull-Compact, and novel densification strategies, to eliminate these bottlenecks.
@@
  • We propose a high-performance framework, LiteGS, to systematically accelerate 3DGS training through system and algorithm codesign. We have open-sourced the framework to the community. [^1] https://github.com/MooreThreads/LiteGS.
@@
  • We perform extensive experiments and ablation studies to evaluate the effectiveness of LiteGS. Compared to the original 3DGS framework, it reduces the training time by up to 13.413.4x while still retaining comparable quality. In high-quality setting, it accelerates the quality SOTA solution by up to 10.810.8x with even slightly better quality.
(b) PSNR versus training time. LiteGS converges to high-quality results over 1010x faster than prior works, demonstrating strong advantages in both compactness and training efficiency.
Figure 1: Reconstruction quality under varying model budgets on the Mip-NeRF 360 garden scene.
2 Related Work
2.1 Acceleration
@@
Although 3DGS is already faster to optimize than classic NeRF, the training still takes tens of minutes or even hours. A key focus in recent 3DGS research has been on removing redundant computations to speed up training. Several works reduce the per-pixel Gaussian overhead by culling or skipping negligible contributions (Liu et al., 2025; Hanson et al., 2025a; Feng et al., 2025). Hou et al. (2025) approximate the alpha blending with a commutative weighted sum, removing the need for per-frame sorting and modestly improving performance on mobile GPUs. Another line of work is pruning redundant primitives. LightGaussian (Fan et al., 2024) proposes a heuristic pruning pipeline. PUP 3D-GS (Hanson et al., 2025b) uses a Hessian-based sensitivity metric to remove  90% of Gaussians from a pretrained model while preserving detail. Other works try to improve the GPU utilizations of 3DGS training. To reduce atomic add operations in the gradient accumulation step, DISTWAR introduced an optimized reduction primitive that leverages warp-level parallelism(Durvasula & Chai, 2024). To remove atomic conflicts on the same tile, Taming 3DGS proposed a ’per-splat parallelization’ approach for backpropagation (Mallick et al., 2024). TC-GS uses Tensor Core units for splatting operations, formulating alpha-blending calculations as matrix multiplies so that otherwise underutilized hardware can accelerate them (Liao et al., 2025).
2.2 Densification Strategies
@@
A key to 3DGS’s success is its ability to refine the point cloud during training via Adaptive Density Control (ADC) - i.e. adding Gaussians in undersampled regions and removing those with negligible opacity. However, the original ADC heuristic has limitations that often produce an explosion of Gaussians with diminishing importance. Recent works have therefore revised the densification logic to be more principled and efficient. Bulò et al. (2024) propose an error-driven densification scheme: instead of using just gradient magnitude, they compute an auxiliary per-pixel error map and spawn new Gaussians where the reconstruction error remains high. Deng et al. (2024) tackle the root cause of Gaussian overshoot by replacing the clone-and-split step with a single long-axis split that avoids overlapping the child Gaussians. Their efficient density control (EDC) method achieves the same or better fidelity using only  1/3 of the Gaussians, by preventing the creation of redundant, overlapping splats. Other strategies guide densification using additional cues. Kim et al. (2024) incorporate color residuals (via spherical harmonics gradients) into the densification criterion, which helps target new Gaussians to areas of color discrepancy rather than just geometric edges. Taming 3DGS (Mallick et al., 2024) uses a score-based selection that adds only the Gaussians which maximally improve the scene’s SSIM/PSNR, and it grows the model to a predetermined budget (rather than indefinitely). By steering new Gaussians to where they raise fidelity the most and stopping when the budget is reached, this method obtains high reconstruction quality with 4-5x fewer Gaussians than the naive ADC (Kerbl et al., 2023).
3 Methodology
@@
where $s(i)$ denotes the index mapping function for depth-ordered primitives in camera space, corresponding to the $i$-th sorted primitive in the Gaussian scene representation $\mathbb{G}$, $T$ is the transmittance $T_{s(i),\pi_{j}}(x,y)=T_{s(i-1),\pi_{j}}(x,y)\cdot(1-\alpha_{s(i-1),\pi_{j}}(x,y))$, $N^{\prime}$ represents the number of primitives visible within the view frustum, $\alpha$ is the opacity term $\alpha_{i,\pi_{j}}(x,y)=o_{i}G(x_{i,\pi_{j}}^{screen}-x,y_{i,\pi_{j}}^{screen}-y,\Sigma_{i,\pi_{j}}^{screen})$ and $G$ is defined as,
$G(\Delta x,\Delta y,\Sigma)=e^{-0.5(\Delta x,\Delta y)\Sigma(\Delta x,\Delta y)^{T}}$ (2)
We use $\mathbb{P}=\{P_{j}|j=1,...,N_{tile}\}$ to represent the $j$-th rendering tiles, and the gradient of $c^{r}_{s(i)}$ is computed by Eq. 3.
$\frac{\partial Loss}{\partial c^{r}_{s(i)}}=\sum_{j=0}^{N_{tile}}\sum_{(x,y)\in P_{j}}\frac{\partial Loss}{\partial I_{r}(x,y)}\cdot T_{s(i),\pi}(x,y)\cdot\alpha_{s(i),\pi}(x,y)$ (3)
3.1 Warp-based Raster
@@
To remarkably eliminate the overhead, we propose a novel warp-based rasterization that only requires a single warp reduction. The key idea is to assign only one warp to each render tile, and let each thread within the warp process multiple pixels through iterative loops. This design first performs an intra-thread reduction operation to accumulate the gradients generated by each pixel within a thread, and then applies a single warp-level reduction to aggregate the gradients for the entire tile. By shrinking the number of expensive warp reductions required per tile to just one, we substantially improve the performance of gradient reduction. The implementation details and a schematic diagram of our Warp-based Rasterization pipeline are provided in Appendix A.1.
However, the design does not come for free. It possibly leads to sharply increased register pressure because the approach tends to assign more pixels to a single thread, which inevitably needs to allocate more temporary variables (such as color and opacity during alpha blending) in the local register. Excessive register occupation causes register spilling, which in turn leads to notable GPU under-utilization. We propose two techniques to ease the register pressure.
3.1.1 Scanline Algorithm
The first technique, inspired by the scanline algorithm(Foley et al., 1996; Shirley et al., 2015) used in software rasterization, reforms the computation of the 2D Gaussian function to maximize data reuse hence diminish register allocation in each thread. This idea lies in the observation that there is a considerable amount of redundant computations during processing multiple pixels for the same Gaussian primitive.
Along any scanline (e.g., the y-axis), the exponent of the Gaussian function (Eq. 2) can be decomposed into a quadratic polynomial of the pixel offset $j$: $\text{Basic}+\text{Linear}\cdot j+\text{Quad}\cdot j^{2}$. This structure allows for significant optimization: the coefficients are computed only once per scanline. The values for all subsequent pixels are then found efficiently through a few incremental multiply-add operations. The derivation of the formulas for the forward and backward phases is shown in Section A.2. The new 2D Gaussian scanline formulation brings two benefits. First, it effectively reduces the variables needed for each pixel calculation from 77 $(x,x^{\text{screen}}_{i},y,y^{\text{screen}}_{i},a,b,c)$ to 33 $(\text{basic},\text{linear},\text{quad})$. Second, it cuts the number of required multiplications (or Fused Multiply and Add) by $G(\Delta x,\Delta y,\Sigma^{\text{screen}})$ from 9 to 2. The same principle applies to the backward pass, where gradients for different pixels with respect to the Basic, Linear, and Quad terms are first summed within the thread, saving both registers and computation.
3.1.2 Strategic Mixed-Precision Computing
@@
These synergistic optimizations help us effectively tune the best performance given the hardware resource constraints. For example, based on the optimized per-pixel register usage, we discovered that each thread processes 4 pixels is the optimal configuration. Using the “one warp per tile” principle, we identified that tile size with $4\times 32=128$ pixels yields the best performance. This size fits the major NVIDIA graphic GPUs since they have the same register file size.
3.2 Cluster-Cull-Compact
@@
To address these challenges, we adapt the established “Cluster-Cull-Compact” paradigm from the field of real-time rendering (Laine et al., 2011; Hapala et al., 2011), tailoring it for the dynamic training process of 3DGS. This pipeline consists of three stages: 1) grouping primitives into spatially contiguous clusters; 2) performing culling at the cluster granularity; and 3) compacting the data of visible clusters into a continuous memory buffer. However, traditional clustering methods like Bounding Volume Hierarchies (BVH) or graph partitioning(Karis, 2021) have high computational complexity for a training loop. The cornerstone of our approach is the utilization of Morton coding (Z-order curve) to achieve high-performance locality-aware sorting(Morton, 1966). By mapping 3D coordinates to a 1D sequence that preserves locality, Morton coding enables us to efficiently enforce spatial coherence at a minimal cost and facilitate lightweight clustering operations.
3.2.1 Morton Codes
Our pipeline begins by normalizing the Gaussian coordinates $p_{i}=(x_{i},y_{i},z_{i})$ into the unit cube $0,1]^{3}$. To generate a 64-bit Morton code, the normalized coordinates are scaled by $2^{21}-1$ and converted to integers. The 64-bit Morton code is then generated by interleaving the bits of the integer coordinates: $\text{Morton}(x,y,z)=\bigoplus_{k=0}^{20}\left(x_{k}\ll(2k)\,|\,y_{k}\ll(2k+1)\,|\,z_{k}\ll(2k+2)\right)$, where $\oplus$ denotes bit concatenation, and $x_{k}$ represents the $k$-th bit of the quantized coordinates.